Archive for January 2017

Apiary Custodians Meeting 22 Jan 2017

Graham, Liz, Rowan and Mel carried out quick check on general status of apiary and colony stores.  Weather bright and sunny but temperature around freezing.  Decision made not to treat with oxalic acid.

Hive #2 – given fondant

Hive #4 – given fondant

Hive #7A – colony dead.  Stores still available in supers so assume not enough bees to form cluster and create heat.   Frames with stores removed and given to #8 and #15F.

Hive #8 – not many bees. Added spare super from #7A.

Hive #15F – added super from 7A.

Combined Nuc – given fondant.

All other hives OK.

Maintenance work to be carried out over the next couple of weeks to clear ferns to the south-west of the apiary (avoiding badger sets!).  Once cleared, Nucs to be sited here.

Next custodians meeting 5 Feb at 1pm.